A scientist begins with 250 grams of a radioactive substance. After 250 minutes, the sample has decayed to 16 grams. Write an exponential equation
f(t)
representing this situation. (Let f be the amount of radioactive substance in grams and t be the time in minutes.)
To the nearest minute what is the half life?
1
Expert's answer
2019-10-29T08:07:07-0400

"f(t)=250e^{-kt}"

After 250 min: "16=250e^{-250k}"

So, "k=\frac{ln250-ln16}{250}=0.011"

To find the half life we should solve: "0.5=e^{-0.011t}"

Thus the half life: "t=\frac{ln0.5}{-0.11}=63 \;min."